# On Closed-Form Expressions for the Fisher–Rao Distance

### v2

@@ 4 Discrete Distributions
Poisson $\lambda^{x}e^{-\lambda}/x!$ $x\in\mathbb{N}$ $\lambda\in\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{1}{\lambda}\end{pmatrix}$ $2\left|\sqrt{\lambda_{1}}-\sqrt{\lambda_{2}}\right|$ [5, 14, 16]
Geometric $p(1-\theta)^{x-1}$ $x\in\mathbb{N}^{*}$ $\theta\in\left]0,1\right[$ $\begin{pmatrix}\frac{1}{\theta^{2}(1-\theta)}\end{pmatrix}$ $2\left|\operatorname{arctanh}\sqrt{1-\theta_{1}}-\operatorname{arctanh}\sqrt{1-\theta_{2}}\right|$ [16, 54]
Negative binomial $\binom{x+r-1}{r-1}\theta^{r}(1-\theta)^{x}$ $x\in\mathbb{N}$ $\theta\in\left]0,1\right[$ $\begin{pmatrix}\frac{r}{\theta^{2}(1-\theta)}\end{pmatrix}$ $2\sqrt{r}\left|\operatorname{arctanh}\sqrt{1-\theta_{1}}-\operatorname{arctanh}\sqrt{1-\theta_{2}}\right|$ [14, 54]
Categorical $\sum_{i=1}^{n}p_{i}\mathbbm{1}_{\{i\}}(x)$ $x\in\{1,\dots,n\}$ $(p_{1},\dots,p_{n-1})\in\left]0,1\right[^{n-1},$ $\sum_{i}p_{i}<1$, $p_{n}\coloneqq 1-\sum_{i}p_{i}$ $\begin{pmatrix}\frac{\delta_{ij}}{p_{i}}+\frac{1}{p_{n}}\end{pmatrix}_{i,j}$ $2\arccos\left(\sum_{i=1}^{n}\sqrt{p_{i}q_{i}}\right)$ [5, 6] [14, 16]
Multinomial $n!\prod_{i=1}^{n}p_{i}^{x_{i}}/x_{i}!$ $(x_{1},\dots,x_{n})\in\mathbb{N}^{n},$ $\sum_{i}x_{i}=m$ $(p_{1},\dots,p_{n-1})\in\left]0,1\right[^{n-1},$ $\sum_{i}p_{i}<1$, $p_{n}\coloneqq 1-\sum_{i}p_{i}$ $\begin{pmatrix}\frac{n\delta_{ij}}{p_{i}}+\frac{n}{p_{n}}\end{pmatrix}_{i,j}$ $2\sqrt{n}\arccos\left(\sum_{i=1}^{n}\sqrt{p_{i}q_{i}}\right)$ [5, 14, 16]
Negative multinomial $p_{n}^{x_{n}}\frac{\Gamma(\sum_{i=1}^{n}x_{i})}{\Gamma(x_{n})}\prod_{i=1}^{n-1}\frac{p_{i}^{x_{i}}}{x_{i}!}$ $(x_{1},\dots,x_{n-1})$         $\in\mathbb{N}^{m-1}$ $(p_{1},\dots,p_{n-1})\in\left]0,1\right[^{n-1},$ $\sum_{i}p_{i}<1$, $p_{n}\coloneqq 1-\sum_{i}p_{i}$ $\begin{pmatrix}\frac{x_{n}\delta_{ij}}{p_{i}p_{n}}+\frac{x_{n}}{p_{n}^{2}}\end{pmatrix}_{i,j}$ $2\sqrt{x_{n}}\operatorname{arccosh}\left(\frac{1-\sum_{i=1}^{n-1}\sqrt{p_{i}q_{i}}}{\sqrt{p_{n}q_{n}}}\right)$ [14, 54]
@@ 4.1 Binomial
A binomial distribution [5, 14, 16] models the probability of having $x$ successes in $n$ i.i.d. Bernoulli experiments with parameter $\theta$. Its p.m.f. is given by $p(x)=\binom{n}{x}\theta^{x}(1-\theta)^{n-x}$, defined for $x\in\left\{0,1,\dots,n\right\}$, and parametrised by $\theta\in\left]0,1\right[$. In this case, $\partial_{\theta}\ell(\theta)=\frac{x}{\theta}-\frac{n-x}{1-\theta}$, and the Fisher information is
@@ 4.2 Poisson
A Poisson distribution [5, 14, 16] has p.m.f. $p(x)=\lambda^{x}e^{-\lambda}/x!$, defined for $x\in\mathbb{N}$, and parametrised by $\lambda\in\mathbb{R}_{+}^{*}$. In this case, we have $\partial_{\lambda}\ell(\lambda)=\frac{x}{\lambda}-1$, and the Fisher information is given by
@@
$=\frac{1}{\lambda^{2}}\mathbb{E}\left[X^{2}\right]-\frac{2}{\lambda}\mathbb{E}\left[X\right]+1$
@@
where we have used that $\mathbb{E}[X]=\lambda$ e $\mathbb{E}[X^{2}]=\lambda(\lambda+1)$. Thus the Fisher–Rao distance is
@@ 4.4 Negative binomial
Consider a sequence of i.i.d. Bernoulli experiments with parameter $\theta$. A negative binomial distribution [14, 54] models the excess of experiments needed until a number of $r$ successes occur. It has p.m.f. $p(x)=\binom{x+r-1}{r-1}\theta^{r}(1-\theta)^{x}$, defined for $x\in\mathbb{N}$, and is parametrised by $\theta\in\left]0,1\right[$, for a fixed222 Since $r\in\mathbb{N}^{*}$ is discrete, it cannot be part of the parametrisation of the statistical manifold, in the differential-geometric sense. For each $r\in\mathbb{N}^{*}$, we have a one-dimensional statistical manifold. $r\in\mathbb{N}^{*}$. We have $\partial_{\theta}\ell(\theta)=\frac{r}{\theta}-\frac{x}{1-\theta}$, and the Fisher information is
@@ 4.5 Categorical
A categorical distribution [5, 6, 14, 16] models a random variable taking values in the sample space $\mathcal{X}=\{1,2,\dots,n\}$ with probabilities $p_{1},\dots,p_{n}$, and has p.m.f. $p(x)=\sum_{i=1}^{n}p_{i}\mathbbm{1}_{\{i\}}(x)$. The associated $(n-1)$-dimensional statistical manifold
@@
The reparametrisation to the sphere provides a nice geometrical interpretation for the relation between the Fisher–Rao distance and the Hellinger distance [69, § 2.4] $d_{\mathrm{H}}(p,q)=\sqrt{\sum_{i=1}^{n}\left(\sqrt{p_{i}}-\sqrt{q_{i}}\right)^{2}}$. While the Fisher–Rao distance between distributions $p$ and $q$ is the length of the radius-two circumference arc between $f(p)$ and $f(q)$, the Hellinger distance is half the Euclidean distance between $f(p)$ and $f(q)$, i.e., $2d_{\mathrm{H}}(p,q)=\|f(p)-f(q)\|_{2}$. In other words, double the Hellinger distance is the arc-chord approximation for the Fisher–Rao distance.
@@ 4.6 Multinomial
Consider $m$ i.i.d. experiments that follow a categorical distribution with $n$ possible outcomes and probabilities $p_{1},\dots,p_{n}$. A multinomial distribution [5, 14, 16] gives the probability of getting $x_{i}$ times the $i$-th outcome, for $1\leq i\leq n$ and $\sum_{i=1}^{n}x_{i}=m$. Its p.m.f. is $p(\boldsymbol{x})=p(x_{1},\dots,x_{n})=n!\prod_{i=1}^{n}p_{i}^{x_{i}}/x_{i}!$ and is defined on the sample space $\mathcal{X}=\left\{\boldsymbol{x}=(x_{1},\dots,x_{n})\in\mathbb{N}^{n}\;\middle|\;\sum_{i=1}^{n}x_{i}=m\right\}$. This distribution is parametrised by the same $\xi=(\xi^{1},\dots,\xi^{n-1})\in\Xi$ as in the categorical distribution, cf. (4.9), with $p_{i}=\xi^{i}$, for $11\leq i\leq n-$ 1 and $p_{n}=1-\sum_{i=1}^{n-1}\xi^{i}$. In this case, we have $\partial_{i}\ell(\xi)=\frac{x_{i}}{p_{i}}-\frac{x_{n}}{p_{n}}$ and $\partial_{j}\partial_{i}\ell(\xi)=-\left(\frac{x_{i}}{p_{i}^{2}}\delta_{ij}+\frac{x_{n}}{p_{n}^{2}}\right)$. Thus, the elements of the Fisher matrix are given by
@@
$=n\left(\frac{\delta_{ij}}{p_{i}}+\frac{1}{p_{n}}\right)$
$=n\left(\frac{\delta_{ij}}{\xi^{i}}+\frac{1}{1-\sum_{k=1}^{n-1}\xi^{k}}\right),$ (4.13)
where we have used that $\mathbb{E}[X_{i}]=np_{i}$. Note that this is the same metric as for the categorical distribution (4.5), up to the factor $n$. Therefore, the Fisher–Rao distance between two distributions $p=\varphi(p_{1},\dots,p_{n-1})$ and $q=\varphi(q_{1},\dots,q_{n-1})$, with $p_{n}=1-\sum_{i=1}^{n-1}p_{i}$ and $q_{n}=1-\sum_{i=1}^{n-1}q_{i}$ is
$d_{\mathrm{FR}}(p,q)=2\sqrt{n}\arccos\left(\sum_{i=1}^{n}\sqrt{p_{i}q_{i}}\right).$ (4.14)
@@ 4.7 Negative multinomial
The negative multinomial distribution [14, 54, 33] generalises the negative binomial. Consider a sequence of i.i.d. categorical experiments with $n$ possible outcomes. A negative multinomial distribution models the number of times $x_{1},\dots,x_{n-1}$ that the first $n-1$ outcomes occur before the $n$-th outcome occurs $x_{n}$ times. It is characterised by the p.m.f. $p(\boldsymbol{x})=p(x_{1},\dots,x_{n-1})=p_{n}^{x_{n}}\frac{\Gamma(\sum_{i=1}^{n}x_{i})}{\Gamma(x_{n})}\prod_{i=1}^{n-1}\frac{p_{i}^{x_{i}}}{x_{i}!}$, and defined for $\boldsymbol{x}=(x_{1},\dots,x_{n-1})\in\mathbb{N}^{n-1}$. It has the same parametrisation as the categorical distribution, cf. (4.9), with $p_{i}=\xi^{i}$, for $11\leq i\leq n-$ 1 and $p_{n}=1-\sum_{i=1}^{n-1}\xi^{i}$, for a fixed $x_{n}\in\mathbb{N}$. In this case, we have $\partial_{i}\ell(\xi)=\frac{x_{i}}{p_{i}}-\frac{x_{n}}{p_{n}}$ and $\partial_{j}\partial_{i}\ell(\xi)=-\frac{x_{i}}{p_{i}^{2}}\delta_{ij}-\frac{x_{n}}{p_{n}^{2}}$. The elements of the Fisher matrix are then
$g_{ij}(\xi)$ $=-\mathbb{E}\left[\partial_{j}\partial_{i}\ell(\xi)\right]=\mathbb{E}\left[-\frac{X_{i}}{p_{i}^{2}}-\frac{x_{n}}{p_{n}^{2}}\right]$
@@
$=\frac{x_{n}}{1-\sum_{i=1}^{n-1}\xi^{i}}\left(\frac{\delta_{ij}}{\xi^{i}}+\frac{1}{1-\sum_{i=1}^{n-1}\xi^{i}}\right),$ (4.15)
@@
$p$ $\mapsto\left(2\sqrt{p_{1}},\dots,2\sqrt{p_{n}}\right).$
@@
$d_{\mathrm{FR}}(p,q)$ $=2\sqrt{x_{n}}d_{\mathcal{H}^{n}}\left((p_{1},\dots,p_{n}),(q_{1},\dots,q_{n})\right)$
@@
It is interesting to note the similar maps (4.11) and (4.16), respectively, for categorical and negative multinomial distributions are used to embed the statistical manifold (in correspondence with the simplex) in the radius-two sphere, but with different metrics. Figure 2 illustrates the geodesics according to these two metrics.
@@ 5 Continuous Distributions
Rayleigh $\frac{x}{\sigma^{2}}\exp\left(-\frac{x^{2}}{2\sigma^{2}}\right)$ $x\in\mathbb{R}_{+}$ $\sigma\in\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{4}{\sigma^{2}}\end{pmatrix}$ $2\left|\log\sigma_{1}-\log\sigma_{2}\right|$**
Erlang $\lambda^{k}x^{k-1}e^{-\lambda x}/(k-1)!$ $x\in\mathbb{R}_{+}$ $\lambda\in\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{k}{\lambda^{2}}\end{pmatrix}$ $\sqrt{k}\left|\log\lambda_{1}-\log\lambda_{2}\right|$**
Gaussian $\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right)$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{1}{\sigma^{2}}&0\\ 0&\frac{2}{\sigma^{2}}\end{pmatrix}$ $2\sqrt{2}\operatorname{arctanh}\left(\sqrt{\frac{(\mu_{1}-\mu_{2})^{2}+2(\sigma_{1}-\sigma_{2})^{2}}{(\mu_{1}-\mu_{2})^{2}+2(\sigma_{1}+\sigma_{2})^{2}}}\right)$ [5, 4, 14] [16, 22]
@@
Generalised Gaussian $\frac{\beta}{2\sigma\Gamma(1/\beta)}\exp\left(-\frac{|x-\mu|^{\beta}}{\sigma}\right)$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\beta}{\sigma^{2}}\frac{\Gamma(2-1/\beta)}{\Gamma(1+1/\beta)}&0\\ 0&\frac{\beta}{\sigma^{2}}\end{pmatrix}$ $\sqrt{\beta+1}\operatorname{arctanh}\left(\sqrt{\frac{(\mu_{1}-\mu_{2})^{2}\Gamma\left(2-\frac{1}{\beta}\right)+\frac{\beta+1}{\beta}(\sigma_{1}-\sigma_{2})^{2}\Gamma\left(1+\frac{1}{\beta}\right)}{(\mu_{1}-\mu_{2})^{2}\Gamma\left(2-\frac{1}{\beta}\right)+\frac{\beta+1}{\beta}(\sigma_{1}+\sigma_{2})^{2}\Gamma\left(1+\frac{1}{\beta}\right)}}\right)$ particular case in [70]
@@
Student’s $t$ $\left(1+\frac{1}{\nu}(\frac{x-\mu}{\sigma})^{2}\right)^{-\frac{\nu+1}{2}}\frac{\Gamma\left((\nu+1)/2\right)}{\sigma\sqrt{\pi\nu}\Gamma(\nu/2)}$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\nu+1}{(\nu+3)\sigma^{2}}&0\\ 0&\frac{2\nu}{(\nu+3)\sigma^{2}}\end{pmatrix}$ $2\sqrt{\frac{2\nu}{\nu+3}}\operatorname{arctanh}\left(\sqrt{\frac{(\nu+1)(\mu_{2}-\mu_{1})^{2}+2\nu(\sigma_{2}-\sigma_{1})^{2}}{(\nu+1)(\mu_{2}-\mu_{1})^{2}+2\nu(\sigma_{2}+\sigma_{1})^{2}}}\right)$ [46]

### v3

@@ 3 Discrete Distributions
Poisson $\frac{\lambda^{x}e^{-\lambda}}{x!}$ $x\in\mathbb{N}$ $\lambda\in\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{1}{\lambda}\end{pmatrix}$ $2\left|\sqrt{\lambda_{1}}-\sqrt{\lambda_{2}}\right|$ [2, 10, 12]
Geometric $\theta(1-\theta)^{x-1}$ $x\in\mathbb{N}^{*}$ $\theta\in\left]0,1\right[$ $\begin{pmatrix}\frac{1}{\theta^{2}(1-\theta)}\end{pmatrix}$ $2\left|\operatorname{arctanh}\sqrt{1-\theta_{1}}-\operatorname{arctanh}\sqrt{1-\theta_{2}}\right|$ [2, 19]
Negative binomial $\frac{\Gamma(x+r)}{x!\Gamma(r)}\theta^{r}(1-\theta)^{x}$ $x\in\mathbb{N}$ $\theta\in\left]0,1\right[$ $\begin{pmatrix}\frac{r}{\theta^{2}(1-\theta)}\end{pmatrix}$ $2\sqrt{r}\left|\operatorname{arctanh}\sqrt{1-\theta_{1}}-\operatorname{arctanh}\sqrt{1-\theta_{2}}\right|$ [12, 19]
Categorical $\sum_{i=1}^{n}p_{i}\mathbbm{1}_{\{i\}}(x)$ $x\in\{1,\dots,n\}$ $(p_{1},\dots,p_{n-1})\in\left]0,1\right[^{n-1},$ $\sum_{i}p_{i}<1$, $p_{n}\coloneqq 1-\sum_{i}p_{i}$ $\begin{pmatrix}\frac{\delta_{ij}}{p_{i}}+\frac{1}{p_{n}}\end{pmatrix}_{i,j}$ $2\arccos\left(\sum_{i=1}^{n}\sqrt{p_{i}q_{i}}\right)$ [2, 4, 10] [12, 46]
Multinomial $m!\prod_{i=1}^{n}\frac{p_{i}^{x_{i}}}{x_{i}!}$ $(x_{1},\dots,x_{n})\in\mathbb{N}^{n},$ $\sum_{i}x_{i}=m$ $(p_{1},\dots,p_{n-1})\in\left]0,1\right[^{n-1},$ $\sum_{i}p_{i}<1$, $p_{n}\coloneqq 1-\sum_{i}p_{i}$ $\begin{pmatrix}\frac{m\delta_{ij}}{p_{i}}+\frac{m}{p_{n}}\end{pmatrix}_{i,j}$ $2\sqrt{m}\arccos\left(\sum_{i=1}^{n}\sqrt{p_{i}q_{i}}\right)$ [2, 10, 12]
Negative multinomial $p_{n}^{x_{n}}\frac{\Gamma(\sum_{i=1}^{n}x_{i})}{\Gamma(x_{n})}\prod_{i=1}^{n-1}\frac{p_{i}^{x_{i}}}{x_{i}!}$ $(x_{1},\dots,x_{n-1})$         $\in\mathbb{N}^{n-1}$ $(p_{1},\dots,p_{n-1})\in\left]0,1\right[^{n-1},$ $\sum_{i}p_{i}<1$, $p_{n}\coloneqq 1-\sum_{i}p_{i}$ $\begin{pmatrix}\frac{x_{n}\delta_{ij}}{p_{i}p_{n}}+\frac{x_{n}}{p_{n}^{2}}\end{pmatrix}_{i,j}$ $2\sqrt{x_{n}}\operatorname{arccosh}\left(\frac{1-\sum_{i=1}^{n-1}\sqrt{p_{i}q_{i}}}{\sqrt{p_{n}q_{n}}}\right)$ [12, 19, 48]
@@ 3.1 Binomial
A binomial distribution [10, 12, 2] models the probability of having $x$ successes in $n$ independent and identically distributed (i.i.d.) Bernoulli experiments with parameter $\theta$. Its p.m.f. is given by $p(x)=\binom{n}{x}\theta^{x}(1-\theta)^{n-x}$, defined for $x\in\{0,1,\dots,n\}$, and parametrised by $\theta\in\left]0,1\right[$. In this case, $\partial_{\theta}\ell(\theta)=\frac{x}{\theta}-\frac{n-x}{1-\theta}$, and the Fisher information is
@@ 3.2 Poisson
A Poisson distribution [10, 12, 2] has p.m.f. $p(x)=\frac{\lambda^{x}e^{-\lambda}}{x!}$, defined for $x\in\mathbb{N}$, and parametrised by $\lambda\in\mathbb{R}_{+}^{*}$. In this case, we have $\partial_{\lambda}\ell(\lambda)=\frac{x}{\lambda}-1$, and the Fisher information is given by
@@
$=\frac{1}{\lambda^{2}}\mathbb{E}[X^{2}]-\frac{2}{\lambda}\mathbb{E}\left[X\right]+1$
@@
where we have used that $\mathbb{E}[X]=\lambda$ and $\mathbb{E}[X^{2}]=\lambda(\lambda+1)$. Thus the Fisher–Rao distance is
@@ 3.4 Negative binomial
A negative binomial distribution [12, 19] models the excess of i.i.d. Bernoulli experiments with parameter $\theta$ needed until a number of $r$ successes occur. It has p.m.f. $p(x)=\frac{\Gamma(x+r)}{x!\Gamma(r)}\theta^{r}(1-\theta)^{x}$, defined for $x\in\mathbb{N}$, and is parametrised by $\theta\in\left]0,1\right[$, for a fixed $r$, that can be extend to $r\in\mathbb{R}^{*}_{+}$. We have $\partial_{\theta}\ell(\theta)=\frac{r}{\theta}-\frac{x}{1-\theta}$, and the Fisher information is
@@ 3.5 Categorical
A categorical distribution [10, 4, 12, 46, 2] models a random variable taking values in the sample space $\mathcal{X}=\{1,2,\dots,n\}$ with probabilities $p_{1},\dots,p_{n}$, and has p.m.f. $p(x)=\sum_{i=1}^{n}p_{i}\mathbbm{1}_{\{i\}}(x)$. The associated $(n-1)$-dimensional statistical manifold
@@
The reparametrisation to the sphere provides a nice geometrical interpretation for the relation between the Fisher–Rao distance and the Hellinger distance [47, § 2.4] $d_{\mathrm{H}}(p,q)=\sqrt{\sum_{i=1}^{n}\left(\sqrt{p_{i}}-\sqrt{q_{i}}\right)^{2}}$. While the Fisher–Rao distance between distributions $p$ and $q$ is the length of the radius-two circumference arc between $f(p)$ and $f(q)$, the Hellinger distance is half the Euclidean distance between $f(p)$ and $f(q)$, i.e., $2d_{\mathrm{H}}(p,q)=\|f(p)-f(q)\|_{2}$. In other words, double the Hellinger distance is the arc-chord approximation for the Fisher–Rao distance.
@@ 3.6 Multinomial
Consider $m$ i.i.d. experiments that follow a categorical distribution with $n$ possible outcomes and probabilities $p_{1},\dots,p_{n}$. A multinomial distribution [10, 12, 2] gives the probability of getting $x_{i}$ times the $i$-th outcome, for $1\leq i\leq n$ and $\sum_{i=1}^{n}x_{i}=m$. Its p.m.f. is $p(\boldsymbol{x})=p(x_{1},\dots,x_{n})=m!\prod_{i=1}^{n}\frac{p_{i}^{x_{i}}}{x_{i}!}$ and is defined on the sample space $\mathcal{X}=\left\{\boldsymbol{x}=(x_{1},\dots,x_{n})\in\mathbb{N}^{n}\;\middle|\;\sum_{i=1}^{n}x_{i}=m\right\}$. This distribution is parametrised by the same $\xi=(\xi^{1},\dots,\xi^{n-1})\in\Xi$ as in the categorical distribution, cf. (3.9), with $p_{i}=\xi^{i}$, for $11\leq i\leq n-$ 1 and $p_{n}=1-\sum_{i=1}^{n-1}\xi^{i}$. In this case, we have $\partial_{i}\ell(\xi)=\frac{x_{i}}{p_{i}}-\frac{x_{n}}{p_{n}}$ and $\partial_{j}\partial_{i}\ell(\xi)=-\frac{x_{i}}{p_{i}^{2}}\delta_{ij}-\frac{x_{n}}{p_{n}^{2}}$. Thus, the elements of the Fisher matrix are given by
@@
$=m\left(\frac{\delta_{ij}}{p_{i}}+\frac{1}{p_{n}}\right)$
$=m\left(\frac{\delta_{ij}}{\xi^{i}}+\frac{1}{1-\sum_{k=1}^{n-1}\xi^{k}}\right),$ (3.13)
where we have used that $\mathbb{E}[X_{i}]=mp_{i}$. Note that this is the same metric as for the categorical distribution (3.5), up to the factor $m$. Therefore, the Fisher–Rao distance between two distributions $p=\varphi(p_{1},\dots,p_{n-1})$ and $q=\varphi(q_{1},\dots,q_{n-1})$, with $p_{n}=1-\sum_{i=1}^{n-1}p_{i}$ and $q_{n}=1-\sum_{i=1}^{n-1}q_{i}$ is
$d_{\mathrm{FR}}(p,q)=2\sqrt{m}\arccos\left(\sum_{i=1}^{n}\sqrt{p_{i}q_{i}}\right).$ (3.14)
@@ 3.7 Negative multinomial
A negative multinomial distribution [12, 19, 48] generalises the negative binomial distribution. Consider a sequence of i.i.d. categorical experiments with $n$ possible outcomes. A negative multinomial distribution models the number of times $x_{1},\dots,x_{n-1}$ that the first $n-1$ outcomes occur before the $n$-th outcome occurs $x_{n}$ times. It is characterised by the p.m.f. $p(\boldsymbol{x})=p(x_{1},\dots,x_{n-1})=p_{n}^{x_{n}}\frac{\Gamma(\sum_{i=1}^{n}x_{i})}{\Gamma(x_{n})}\prod_{i=1}^{n-1}\frac{p_{i}^{x_{i}}}{x_{i}!}$, and defined for $\boldsymbol{x}=(x_{1},\dots,x_{n-1})\in\mathbb{N}^{n-1}$. It has the same parametrisation as the categorical distribution, cf. (3.9), with $p_{i}=\xi^{i}$, for $11\leq i\leq n-$ 1 and $p_{n}=1-\sum_{i=1}^{n-1}\xi^{i}$, for a fixed $x_{n}$ that can be extended to $x_{n}\in\mathbb{R}_{+}^{*}$. In this case, we have $\partial_{i}\ell(\xi)=\frac{x_{i}}{p_{i}}-\frac{x_{n}}{p_{n}}$ and $\partial_{j}\partial_{i}\ell(\xi)=-\frac{x_{i}}{p_{i}^{2}}\delta_{ij}-\frac{x_{n}}{p_{n}^{2}}$. The elements of the Fisher matrix are then
@@
$=\frac{x_{n}}{1-\sum_{k=1}^{n-1}\xi^{k}}\left(\frac{\delta_{ij}}{\xi^{i}}+\frac{1}{1-\sum_{k=1}^{n-1}\xi^{k}}\right),$ (3.15)
@@
$p=\sum_{i=1}^{n}p_{i}\mathbbm{1}_{\{i\}}$ $\mapsto\left(2\sqrt{p_{1}},\dots,2\sqrt{p_{n}}\right).$
@@
$d_{\mathrm{FR}}(p,q)$ $=2\sqrt{x_{n}}\ \widetilde{d}_{S_{2}^{n-1}}\left((2\sqrt{p_{1}},\dots,2\sqrt{p_{n}}),(2\sqrt{q_{1}},\dots,2\sqrt{q_{n}})\right)$
@@
It is interesting to note that the similar maps (3.11) and (3.16), respectively, for categorical and negative multinomial distributions are used to embed the statistical manifold (in correspondence with the simplex) in the radius-two sphere, but with different metrics. Figure 2 illustrates the geodesics according to these two metrics.
@@ 4 Continuous Distributions
Rayleigh $\frac{x}{\sigma^{2}}\exp\left(-\frac{x^{2}}{2\sigma^{2}}\right)$ $x\in\mathbb{R}_{+}$ $\sigma\in\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{4}{\sigma^{2}}\end{pmatrix}$ $2\left|\log\sigma_{1}-\log\sigma_{2}\right|$
Erlang $\lambda^{k}x^{k-1}e^{-\lambda x}/(k-1)!$ $x\in\mathbb{R}_{+}$ $\lambda\in\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{k}{\lambda^{2}}\end{pmatrix}$ $\sqrt{k}\left|\log\lambda_{1}-\log\lambda_{2}\right|$
Gaussian $\frac{1}{\sigma\sqrt{2\pi}}\exp\left(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\right)$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{1}{\sigma^{2}}&0\\ 0&\frac{2}{\sigma^{2}}\end{pmatrix}$ $2\sqrt{2}\operatorname{arctanh}\left(\sqrt{\frac{(\mu_{1}-\mu_{2})^{2}+2(\sigma_{1}-\sigma_{2})^{2}}{(\mu_{1}-\mu_{2})^{2}+2(\sigma_{1}+\sigma_{2})^{2}}}\right)$ [2, 10, 12] [17, 34, 45]
@@
Generalised Gaussian $\frac{\beta}{2\sigma\Gamma(1/\beta)}\exp\left(-\frac{|x-\mu|^{\beta}}{\sigma}\right)$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\beta}{\sigma^{2}}\frac{\Gamma(2-\frac{1}{\beta})}{\Gamma(1+\frac{1}{\beta})}&0\\ 0&\frac{\beta}{\sigma^{2}}\end{pmatrix}$ $\sqrt{\beta+1}\operatorname{arctanh}\left(\sqrt{\frac{(\mu_{1}-\mu_{2})^{2}\Gamma\left(2-\frac{1}{\beta}\right)+\frac{\beta+1}{\beta}(\sigma_{1}-\sigma_{2})^{2}\Gamma\left(1+\frac{1}{\beta}\right)}{(\mu_{1}-\mu_{2})^{2}\Gamma\left(2-\frac{1}{\beta}\right)+\frac{\beta+1}{\beta}(\sigma_{1}+\sigma_{2})^{2}\Gamma\left(1+\frac{1}{\beta}\right)}}\right)$ Particular case in [22]
@@
Student’s $t$ a $\left(1+\frac{1}{\nu}(\frac{x-\mu}{\sigma})^{2}\right)^{-\frac{\nu+1}{2}}$$\times\frac{\Gamma\left((\nu+1)/2\right)}{\sigma\sqrt{\pi\nu}\Gamma(\nu/2)}$ $x\in\mathbb{R}$ $(\mu,\sigma)\in\mathbb{R}\times\mathbb{R}_{+}^{*}$ $\begin{pmatrix}\frac{\nu+1}{(\nu+3)\sigma^{2}}&0\\ 0&\frac{2\nu}{(\nu+3)\sigma^{2}}\end{pmatrix}$ $2\sqrt{\frac{2\nu}{\nu+3}}\operatorname{arctanh}\left(\sqrt{\frac{(\nu+1)(\mu_{2}-\mu_{1})^{2}+2\nu(\sigma_{2}-\sigma_{1})^{2}}{(\nu+1)(\mu_{2}-\mu_{1})^{2}+2\nu(\sigma_{2}+\sigma_{1})^{2}}}\right)$ [17]